# A uniform kernel trick for high and infinite-dimensional two-sample problems

### v2

@@ 3 Main results
In this section we first show that (unions of) unit balls of RKHS are universal Donsker under mild conditions. This is an important technical result of independent interest that is the starting point in the proofs of the asymptotic results. We analyze the asymptotic behaviour of the plug-in estimator (12) of the supremum kernel distance in (4) and (11). The results are quite general as $\operatorname{P}$ and $\operatorname{Q}$ are assumed to be Borel probability measures on a separable metric space. The proofs are based on empirical processes theory together with the (extended) delta method (Shapiro, 1991, Theorem 2.1) and some recent differentiability results for the supremum (Cárcamo et al. (2020)). This differential approach differs from previous methods (as those in Balakrishnan et al. (2012) or Borgwardt et al. (2012)) in which the theory of U-statistics is used to derive the asymptotic results. Our approach has some advantages: it is applicable to variables taking values in general spaces, including functional spaces, and the equal sample size constraint of previous works is removed. Furthermore, the results are applicable in other contexts (such as tests for equality between two copulas) by just changing the underlying stochastic process in the spirit of Cárcamo et al. (2020). Another essential difference between our methodology and other approaches is the way in which the tuning parameter $\lambda$ is treated. The asymptotic theory in Borgwardt et al. (2012) (and other related works) is derived for a fixed kernel, while the experiments incorporate the Gaussian kernel in (2) with a data-driven choice of $\lambda$. As pointed out by the authors, an automatic method for selecting $\lambda$ is an interesting area of research with some theoretical implications: setting the kernel using the sample being tested may cause changes to the asymptotic distribution. Regarding this, we note that our procedure to deal with the tuning parameter $\lambda$ is fully incorporated in the asymptotic analysis thanks to the use of the supremum kernel distance (4).
@@
We list some assumptions for later reference. We briefly explain the meaning and implications of each of them. In what follows, $k$ is a kernel, $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$ a family of kernels (which might come from different parametric families), and $\operatorname{P}$ and $\operatorname{Q}\in\mathcal{M}_{\operatorname{p}}(\mathcal{X})$, Borel probability measures defined on a space $\mathcal{X}$. In what follows we use the standard notation in functional analysis and operator theory; for $k_{1}$ and $k_{2}$ positive definite kernels on $\mathcal{X}$, we denote $k_{1}\ll k_{2}$ if and only if $k_{2}-k_{1}$ is a positive definite kernel; see (Aronszajn, 1950, Part I.7).
@@
Assumptions 3 and 3 together have important consequences. Firstly, they imply that $\mathcal{H}_{k,\lambda}$ is constituted by continuous and bounded functions (see (Berlinet and Thomas-Agnan, 2011, Theorem 17)), therefore measurable and integrable. Moreover, under these two conditions the mean embedding $\mu_{\operatorname{P}}^{\lambda}$ exits (for each $\operatorname{P}$ and $\lambda$). In particular, the supremum kernel distance (4) is well-defined. 3 and 3 are also essential to show that the class $\mathcal{F}_{k,\Lambda}$ in (11) is universal Donsker, which is a key point in the proofs of the following theorems.
Assumption 3 entails that $d_{k,\Lambda}(\operatorname{P},\operatorname{Q})>0$, whenever $\operatorname{P}\neq\operatorname{Q}$, i.e., the supremum kernel distance separates different probability measures. Therefore, $d_{k,\Lambda}$ in (3) is a proper metric on $\mathcal{M}_{\operatorname{p}}(\mathcal{X})$. Regarding this, we recall that a reproducing kernel $k$ is said to be characteristic whenever $d_{k}(\operatorname{P},\operatorname{Q})=0$ if and only if $\operatorname{P}=\operatorname{Q}$, for all $\operatorname{P},\operatorname{Q}\in\mathcal{M}_{\operatorname{p}}(\mathcal{X})$; see Fukumizu et al. (2008). This is equivalent to “integrally strictly positive definiteness”, see (Fukumizu et al., 2010, Theorem 7). Hence, 3 could be understood as the family $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$ being characteristic in the sense that for each pair of different measures there is a kernel in the family separating them. Observe that this condition is necessary to carry out the test $\operatorname{H}_{0}:\operatorname{P}=\operatorname{Q}$ by means of the statistic (12). Otherwise, the two-sample test that we propose only checks the weaker hypothesis $d_{k,\Lambda}(\operatorname{P},\operatorname{Q})=0$. Note that 3 is less demanding than asking for a specific kernel to be characteristic, a standard requirement on this topic. In infinite dimension, necessary and sufficient conditions for the Gaussian kernel to be characteristic are given in Duncan and Wynne (2022). We also observe that 3 is not specifically required to obtain the asymptotic distribution under $\operatorname{H}_{0}$ in Theorem 2.
Finally, 3 is a technical requirement to derive the asymptotic distribution of the test statistic under the alternative hypothesis using the results in Cárcamo et al. (2020). 3 is necessary for the combination of the associated empirical processes to converge.
@@
The hypotheses above can be verified for most families of kernels that are used in practice by properly choosing the parameter space. The most demanding assumption about the kernel family is perhaps 3. In particular, this always ensures the applicability of the results, both in high and infinite dimensions, since the practical implementation of the procedure is carried out by choosing a grid of points in the parameter space; see Section 4. When $\mathcal{X}=\mathbb{R}^{d}$, a finite-dimensional space, the usual parametric families of kernels often generate a nested collection of RKHS; see Zhang and Zhao (2013). This means that for $\lambda_{1},\lambda_{2}\in\Lambda$, there exists a constant $c=c\left(\lambda_{1},\lambda_{2},d\right)$ such that $k_{\lambda_{1}}\ll c\,k_{\lambda_{2}}$ (or the other way around). In such cases, 3 is valid for a compact subset $\Lambda$ of the whole parametric space by using one of the kernels of the family as the bounding kernel $k$ in 3. Some important examples included in this setting are the families of Gaussian and Laplacian kernels, inverse multiquadrics kernels, B-spline kernels, Matérn kernels, among others; see Sriperumbudur (2016) and (Zhang and Zhao, 2013, Theorems 3.5, 3.6, and 3.7).
Nevertheless, the problem is more delicate in infinite dimension. If $\mathcal{X}=\mathbb{R}^{d}$ and for the usual parametric families of kernels, the best constant $c$ in “inequalities” of the form $k_{\lambda_{1}}\ll c\,k_{\lambda_{2}}$ depends on the dimension $d$ and blows up when $d$ goes to infinity; see (Zhang and Zhao, 2013, Theorems 3.5 and 3.6). Therefore, when the domain is functional (for instance, if $\mathcal{X}=\operatorname{L}^{2}([0,1])$), the task of finding a dominating kernel is more involved. An example can be built when the parameter space
@@
is the unit ball of the Cameron-Martin space associated to the Wiener measure in $\mathcal{C}([0,1])$, the space of continuous functions on $[0,1]$. A family of kernels $k_{m}(x,y)$ fulfilling 3 and 3 can be constructed using Minlos-Sazanov Theorem (see e.g., (Duncan and Wynne, 2022, Th. 24)) and relying on ideas by (Wendland, 2004, Chapter 10) and (Zhang and Zhao, 2013, Prop. 3.1). Additionally, we observe that 3 is fulfilled for families of positive linear (or convex) combinations of a finite family of kernels. In this example, the set of parameter $\Lambda$ is given by the weights of the considered combinations; see Balakrishnan et al. (2012). We finally refer to (Berlinet and Thomas-Agnan, 2011, Chapter 7) and (Paulsen and Raghupathi, 2016, Chapter 4) for a wider catalog of families of kernels within this context.
@@ A Donsker property for units balls in RKHS
Establishing that a class of functions is (uniform) Donsker has important consequences. This property is equivalent to having an empirical central limit theorem, which is at the heart of most asymptotic results in statistics. Therefore, this kind of Donsker-type results are relevant by themselves and of independent interest. For example, in (Sriperumbudur, 2016, Theorem 4.3) (see also Giné and Nickl (2008), Giné and Nickl (2016)) it is shown that $\mathcal{F}_{k,\Lambda}$ in (11) is Donsker for some specific finite-dimensional parametric families and for a suitable subset of $\Lambda$. Then, this result is applied to derive asymptotic distributions of kernel density estimators. In Sriperumbudur (2016), the proofs of the Donsker property for RKHS unit balls are obtained when $\mathcal{X}=\mathbb{R}^{d}$ by direct covering (entropy-based) arguments. The underlying bounds in these references depend on the dimension $d$. Therefore, it seems difficult to extend these Donsker-type statements to the infinite-dimensional case. However, Theorem 1 below is suitable for the general framework where $\mathcal{X}$ might be an infinite-dimensional space, and thus useful in statistical problems with functional data. The following theorem establishes that unit balls (and even the union of units balls) of RKHS are universal Donsker. In the first part of the proof (in Section 6) we use (Marcus, 1985, Theorem 1.1), while in the second one we show that the union of unit balls is included in a ball of the space $\mathcal{H}_{k}$ by using Aronszajn’s inclusion theorem ((Aronszajn, 1950, Theorem I)).
@@
This theorem extends (Sriperumbudur, 2016, Theorem 4.3), where the Donsker property was shown under more demanding analytical conditions, to any family of kernels satisfying 3.
@@
where $d_{k,\Lambda}$ is defined in (11), $Z_{j,\lambda}=\left\langle\mathbb{G}_{\operatorname{P}},\varphi_{j,\lambda}\right\rangle_{\mathcal{H}_{k,\lambda}^{\ast}}$ (for each $\lambda\in\Lambda$ and $j\in\mathbb{N}$) and $\varphi_{j,\lambda}$ is the $j$-th eigenfunction of the covariance operator of $\mathbb{G}_{\operatorname{P}}$ on $\mathcal{H}_{k,\lambda}^{\ast}$. Moreover, $\left\{Z_{j,\lambda}\right\}_{j\in\mathbb{N},\lambda\in\Lambda}$ are jointly Gaussian and for a fixed $\lambda\in\Lambda$, $\left\{Z_{j,\lambda}\right\}_{j\in\mathbb{N}}$ are independent with $Z_{j,\lambda}\sim\mathcal{N}\left(0,\beta_{j,\lambda}\right)$, where $\beta_{j,\lambda}$ is the eigenvalue associated to $\varphi_{j,\lambda}$.
In the first step of the proof of this theorem we use Theorem 1 to derive the weak convergence of the underlying process. The rest of the proof is rather technical. The basic ideas are as follows: we use of the continuous mapping theorem to obtain the convergence of the statistic; subsequently, we apply a functional Karhunen-Loève-type theorem in the dual space $\mathcal{H}_{k,\lambda}^{\ast}$ (Lemma 3 in Section 6) to the resulting limiting process to achieve (13). Note that in the family $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$ we can include kernels from different parametric families or mixtures of kernels from distinct families in order to robustify the test statistic. Theorem 2 complements in several directions other previous works on this topic, starting from (Borgwardt et al., 2006, Th. 8). See also Guo et al. (2022), Smaga and Zhang (2022) for more recent references.
@@
The following theorem establishes the asymptotic distribution of (the normalized version) of (12) under the alternative hypothesis of the homogeneity test. Therefore, it provides the consistency of the testing procedure based on the supremum kernel distance. Additionally, this result might be potentially useful in order to develop tests of almost homogeneity, that is, problems in which we are interested in testing $\operatorname{H}_{0}:d_{k,\Lambda}(\operatorname{P},\operatorname{Q})\leq\varepsilon$ versus $\text{H}_{1}:d_{k,\Lambda}(\operatorname{P},\operatorname{Q})>\varepsilon$, for some $\varepsilon>0$. Analogously, this idea is also applicable to provide statistical evidence in favour of almost homogeneity when $\operatorname{H}_{0}$ and $\operatorname{H}_{1}$ above are interchanged. Related ideas can be found in del Barrio et al. (2020) and Dette and Kokot (2022).
@@
Theorem 3 directly provides the consistency of the homogeneity test based on the supremum kernel distance $d_{k,\Lambda}$ in (4). We also observe that $\mathbb{G}$ is a zero mean Gaussian process indexed by $\mathcal{F}_{k,\Lambda}$. Further, $h^{+,\lambda}$ is called witness function in Borgwardt et al. (2006) as the maximum mean discrepancy over $\mathcal{F}_{k,\lambda}$ is attained at this element, that is, $\operatorname{P}\left(h^{+,\lambda}\right)-\operatorname{Q}\left(h^{+,\lambda}\right)=\|\mu_{\operatorname{P}}^{\lambda}-\mu_{\operatorname{Q}}^{\lambda}\|_{\mathcal{H}_{k,\lambda}}$. Therefore, the limit in (14) corresponds to the supremum of $\mathbb{G}$ over the set of witness functions for which the value of the uniform kernel distance is achieved. Regarding the proof of Theorem 3, we mention that the extended delta method (see (Shapiro, 1991, Theorem 2.1)) plays a key role. First, we use Theorem 1 to show that $\mathbb{G}$ is the limit of the underlying process. Later, we adapt some ideas from Cárcamo et al. (2020) to derive (14). The following result is a direct consequence of Theorem 3 when the family of kernels has a single element, $k$.
@@
Corollary 1 extends some previous results in which it is assumed that $n=m$; see (Borgwardt et al., 2006, Th. 8), (Borgwardt et al., 2006a, Th. 2.5), and (Duncan and Wynne, 2022, Th. 16).

### v3

@@ 3 Main results
In this section we first show that (unions of) unit balls of RKHS are universal Donsker under mild conditions. This is an important technical result of independent interest that is the starting point in the proofs of the asymptotic results. We analyze the asymptotic behaviour of the plug-in estimator (12) of the supremum kernel distance in (4) and (11). The results are quite general as $\operatorname{P}$ and $\operatorname{Q}$ are assumed to be Borel probability measures on a separable metric space. The proofs are based on empirical processes theory together with the (extended) delta method [44, Theorem 2.1] and some recent differentiability results for the supremum ([14]). This differential approach differs from previous methods (as those in [3] or [12]) in which the theory of U-statistics is used to derive the asymptotic results. Our approach has some advantages: it is applicable to variables taking values in general spaces, including functional spaces, and the equal sample size constraint of previous works is removed. Furthermore, the results are applicable in other contexts (such as tests for equality between two copulas) by just changing the underlying stochastic process in the spirit of [14].
Another essential difference between our methodology and other approaches is the way in which the tuning parameter $\lambda$ is treated. The asymptotic theory in [12] (and other related works) is derived for a fixed kernel, while the experiments incorporate the Gaussian kernel in (2) with a data-driven choice of $\lambda$. As pointed out by the authors, an automatic method for selecting $\lambda$ is an interesting area of research with some theoretical implications: setting the kernel using the sample being tested may cause changes to the asymptotic distribution. Regarding this, we note that our procedure to deal with the tuning parameter $\lambda$ is fully incorporated in the asymptotic analysis thanks to the use of the supremum kernel distance (4).
@@
We list some assumptions for later reference. We briefly explain the meaning and implications of each of them. In what follows, $k$ is a kernel, $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$ a family of kernels (which might come from different parametric families), and $\operatorname{P}$ and $\operatorname{Q}\in\mathcal{M}_{\operatorname{p}}(\mathcal{X})$, Borel probability measures defined on a space $\mathcal{X}$. In what follows we use the standard notation in functional analysis and operator theory; for $k_{1}$ and $k_{2}$ positive definite kernels on $\mathcal{X}$, we denote $k_{1}\ll k_{2}$ if and only if $k_{2}-k_{1}$ is a positive definite kernel; see [2, Part I.7].
@@
Assumptions 3 and 3 together have important consequences. Firstly, they imply that $\mathcal{H}_{k,\lambda}$ is constituted by continuous and bounded functions (see [6, Theorem 17]), therefore measurable and integrable. Moreover, under these two conditions the mean embedding $\mu_{\operatorname{P}}^{\lambda}$ exits (for each $\operatorname{P}$ and $\lambda$). In particular, the supremum kernel distance (4) is well-defined. 3 and 3 are also essential to show that the class $\mathcal{F}_{k,\Lambda}$ in (11) is universal Donsker, which is a key point in the proofs of the following theorems.
Assumption 3 entails that $d_{k,\Lambda}(\operatorname{P},\operatorname{Q})>0$, whenever $\operatorname{P}\neq\operatorname{Q}$, i.e., the supremum kernel distance separates different probability measures. Therefore, $d_{k,\Lambda}$ in (3) is a proper metric on $\mathcal{M}_{\operatorname{p}}(\mathcal{X})$. Regarding this, we recall that a reproducing kernel $k$ is said to be characteristic whenever $d_{k}(\operatorname{P},\operatorname{Q})=0$ if and only if $\operatorname{P}=\operatorname{Q}$, for all $\operatorname{P},\operatorname{Q}\in\mathcal{M}_{\operatorname{p}}(\mathcal{X})$; see [24]. This is equivalent to “integrally strictly positive definiteness”, see [26, Theorem 7]. Hence, 3 could be understood as the family $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$ being characteristic in the sense that for each pair of different measures there is a kernel in the family separating them. Observe that this condition is necessary to carry out the test $\operatorname{H}_{0}:\operatorname{P}=\operatorname{Q}$ by means of the statistic (12). Otherwise, the two-sample test that we propose only checks the weaker hypothesis $d_{k,\Lambda}(\operatorname{P},\operatorname{Q})=0$. Note that 3 is less demanding than asking for a specific kernel to be characteristic, a standard requirement on this topic. In infinite dimension, necessary and sufficient conditions for the Gaussian kernel to be characteristic are given in [19]. We also observe that 3 is not specifically required to obtain the asymptotic distribution under $\operatorname{H}_{0}$ in Theorem 2.
Finally, 3 is a technical requirement to derive the asymptotic distribution of the test statistic under the alternative hypothesis using the results in [14]. 3 is necessary for the combination of the associated empirical processes to converge.
@@
The hypotheses above can be verified for most families of kernels that are used in practice by properly choosing the parameter space. The most demanding assumption about the kernel family is perhaps 3. In particular, this always ensures the applicability of the results, both in high and infinite dimensions, since the practical implementation of the procedure is carried out by choosing a grid of points in the parameter space; see Section 4.
When $\mathcal{X}=\mathbb{R}^{d}$, a finite-dimensional space, the usual parametric families of kernels often generate a nested collection of RKHS; see [51]. This means that for $\lambda_{1},\lambda_{2}\in\Lambda$, there exists a constant $c=c\left(\lambda_{1},\lambda_{2},d\right)$ such that $k_{\lambda_{1}}\ll c\,k_{\lambda_{2}}$ (or the other way around). In such cases, 3 is valid for a compact subset $\Lambda$ of the whole parametric space by using one of the kernels of the family as the bounding kernel $k$ in 3. Some important examples included in this setting are the families of Gaussian and Laplacian kernels, inverse multiquadrics kernels, B-spline kernels, Matérn kernels, among others; see [46] and [51, Theorems 3.5, 3.6, and 3.7].
Nevertheless, the problem is more delicate in infinite dimension. If $\mathcal{X}=\mathbb{R}^{d}$ and for the usual parametric families of kernels, the best constant $c$ in “inequalities” of the form $k_{\lambda_{1}}\ll c\,k_{\lambda_{2}}$ depends on the dimension $d$ and blows up when $d$ goes to infinity; see [51, Theorems 3.5 and 3.6]. Therefore, when the domain is functional (for instance, if $\mathcal{X}=\operatorname{L}^{2}([0,1])$), the task of finding a dominating kernel is more involved. An example can be built when the parameter space
@@
is the unit ball of the Cameron-Martin space associated to the Wiener measure in $\mathcal{C}([0,1])$, the space of continuous functions on $[0,1]$. A family of kernels $k_{m}(x,y)$ fulfilling 3 and 3 can be constructed using Minlos-Sazanov Theorem (see e.g., [19, Th. 24]) and relying on ideas by [50, Chapter 10] and [51, Prop. 3.1].
Additionally, we observe that 3 is fulfilled for families of positive linear (or convex) combinations of a finite family of kernels. In this example, the set of parameter $\Lambda$ is given by the weights of the considered combinations; see [3]. We finally refer to [6, Chapter 7] and [40, Chapter 4] for a wider catalog of families of kernels within this context.
@@ A Donsker property for units balls in RKHS
Establishing that a class of functions is (uniform) Donsker has important consequences. This property is equivalent to having an empirical central limit theorem, which is at the heart of most asymptotic results in statistics. Therefore, this kind of Donsker-type results are relevant by themselves and of independent interest. For example, in [46, Theorem 4.3] (see also [30], [31]) it is shown that $\mathcal{F}_{k,\Lambda}$ in (11) is Donsker for some specific finite-dimensional parametric families and for a suitable subset of $\Lambda$. Then, this result is applied to derive asymptotic distributions of kernel density estimators. In [46], the proofs of the Donsker property for RKHS unit balls are obtained when $\mathcal{X}=\mathbb{R}^{d}$ by direct covering (entropy-based) arguments. The underlying bounds in these references depend on the dimension $d$. Therefore, it seems difficult to extend these Donsker-type statements to the infinite-dimensional case. However, Theorem 1 below is suitable for the general framework where $\mathcal{X}$ might be an infinite-dimensional space, and thus useful in statistical problems with functional data.
The following theorem establishes that unit balls (and even the union of units balls) of RKHS are universal Donsker. In the first part of the proof (in Section 6) we use [38, Theorem 1.1], while in the second one we show that the union of unit balls is included in a ball of the space $\mathcal{H}_{k}$ by using Aronszajn’s inclusion theorem ([2, Theorem I]).
@@
This theorem extends [46, Theorem 4.3], where the Donsker property was shown under more demanding analytical conditions, to any family of kernels satisfying 3.
@@
where $d_{k,\Lambda}$ is defined in (11), $Z_{j,\lambda}=\left\langle\mathbb{G}_{\operatorname{P}},\varphi_{j,\lambda}\right\rangle_{\mathcal{H}_{k,\lambda}^{\ast}}$ (for each $\lambda\in\Lambda$ and $j\in\mathbb{N}$) and $\varphi_{j,\lambda}$ is the $j$-th eigenfunction of the covariance operator of $\mathbb{G}_{\operatorname{P}}$ on $\mathcal{H}_{k,\lambda}^{\ast}$.
Moreover, $\left\{Z_{j,\lambda}\right\}_{j\in\mathbb{N},\lambda\in\Lambda}$ are jointly Gaussian and for a fixed $\lambda\in\Lambda$, $\left\{Z_{j,\lambda}\right\}_{j\in\mathbb{N}}$ are independent with $Z_{j,\lambda}\sim\mathcal{N}\left(0,\beta_{j,\lambda}\right)$, where $\beta_{j,\lambda}$ is the eigenvalue associated to $\varphi_{j,\lambda}$.
In the first step of the proof of this theorem we use Theorem 1 to derive the weak convergence of the underlying process. The rest of the proof is rather technical. The basic ideas are as follows: we use of the continuous mapping theorem to obtain the convergence of the statistic; subsequently, we apply a functional Karhunen-Loève-type theorem in the dual space $\mathcal{H}_{k,\lambda}^{\ast}$ (Lemma 3 in Section 6) to the resulting limiting process to achieve (13). Note that in the family $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$ we can include kernels from different parametric families or mixtures of kernels from distinct families in order to robustify the test statistic.
Theorem 2 complements in several directions other previous works on this topic, starting from [10, Th. 8]. See also [33], [45] for more recent references.
@@
The following theorem establishes the asymptotic distribution of (the normalized version) of (12) under the alternative hypothesis of the homogeneity test. Therefore, it provides the consistency of the testing procedure based on the supremum kernel distance. Additionally, this result might be potentially useful in order to develop tests of almost homogeneity, that is, problems in which we are interested in testing $\operatorname{H}_{0}:d_{k,\Lambda}(\operatorname{P},\operatorname{Q})\leq\varepsilon$ versus $\text{H}_{1}:d_{k,\Lambda}(\operatorname{P},\operatorname{Q})>\varepsilon$, for some $\varepsilon>0$. Analogously, this idea is also applicable to provide statistical evidence in favour of almost homogeneity when $\operatorname{H}_{0}$ and $\operatorname{H}_{1}$ above are interchanged. Related ideas can be found in [4] and [18].
@@
Theorem 3 directly provides the consistency of the homogeneity test based on the supremum kernel distance $d_{k,\Lambda}$ in (4). We also observe that $\mathbb{G}$ is a zero mean Gaussian process indexed by $\mathcal{F}_{k,\Lambda}$. Further, $h^{+,\lambda}$ is called witness function in [10] as the maximum mean discrepancy over $\mathcal{F}_{k,\lambda}$ is attained at this element, that is, $\operatorname{P}\left(h^{+,\lambda}\right)-\operatorname{Q}\left(h^{+,\lambda}\right)=\|\mu_{\operatorname{P}}^{\lambda}-\mu_{\operatorname{Q}}^{\lambda}\|_{\mathcal{H}_{k,\lambda}}$. Therefore, the limit in (14) corresponds to the supremum of $\mathbb{G}$ over the set of witness functions for which the value of the uniform kernel distance is achieved. Regarding the proof of Theorem 3, we mention that the extended delta method (see [44, Theorem 2.1]) plays a key role. First, we use Theorem 1 to show that $\mathbb{G}$ is the limit of the underlying process. Later, we adapt some ideas from [14] to derive (14).
The following result is a direct consequence of Theorem 3 when the family of kernels has a single element, $k$.
@@
Corollary 1 extends some previous results in which it is assumed that $n=m$; see [10, Th. 8], [11, Th. 2.5], and [19, Th. 16].